OpenStudy (anonymous):

1. Kevin ran 4 miles more than Steve ran. The sum of their distances is 26 miles. How far did Steve run? The domain of the solution is . (a) Define your variable. (b) Write algebraic expressions for the distances each person ran. (c) Write an equation that can help you solve the problem. (d) Solve the equation. (e) Check your answer.

OpenStudy (denonakavro):

Hey tell me what you think about this If Steve ran x and then Kevin ran x+4 So x+x+4=26

OpenStudy (denonakavro):

Lets see a) x would be the distance steve ran and x+4 would be distance Kevin ran b) for Kevin : x+4=26-x for Steve : x= 26-x-4 c) An equation to solve would be x+x+4=26 d) x+x+4 =26 2x+4=26 2x=22 x=11 distance Steve ran x+4 =15 distance Kevin ran e) 11+15 = 26 15-11 =4

OpenStudy (denonakavro):

Hey can you medal and fan me if you think I am correct